OpenStudy (anonymous):

How do you set up the function to this question. Jacob invests $2500. He would earn 95% compounded continuously for 10 years.

OpenStudy (abmon98):

2500*95/100=2375*10=23750+2500=26250

OpenStudy (whpalmer4):

@abmon98 that is not correct for continuous compounding